# Artificial Liver Classifier: A New Alternative to Conventional Machine Learning Models

### v1

@@ Abstract
Supervised machine learning classifiers often encounter challenges related to performance, accuracy, and overfitting. This paper introduces the Artificial Liver Classifier (ALC), a novel supervised learning classifier inspired by the human liver’s detoxification function. The ALC is characterized by its simplicity, speed, hyperparameters-free, ability to reduce overfitting, and effectiveness in addressing multi-classification problems through straightforward mathematical operations. To optimize the ALC’s parameters, an improved FOX optimization algorithm (IFOX) is employed as the training method. The proposed ALC was evaluated on five benchmark machine learning datasets: Iris Flower, Breast Cancer Wisconsin, Wine, Voice Gender, and MNIST. The results demonstrated competitive performance, with the ALC achieving 100% accuracy on the Iris dataset, surpassing logistic regression, multilayer perceptron, and support vector machine. Similarly, on the Breast Cancer dataset, it achieved 99.12% accuracy, outperforming XGBoost and logistic regression. Across all datasets, the ALC consistently exhibited lower overfitting gaps and loss compared to conventional classifiers. These findings highlight the potential of leveraging biological process simulations to develop efficient machine learning models and open new avenues for innovation in the field.
@@ 1 Introduction
In the early stages of AI, researchers focused on building systems (with minimal intelligence) capable of performing specific tasks using fixed rules (conditional and logical operations). As the field evolved, scientists realized that intelligent systems needed methods to learn from data, rather than relying on rigid rule-based methods with minimal capabilities [13, 14]. As a result, supervised learning algorithms, specifically classifiers, emerged as tools for learning systems to make predictions or decisions based on the available experiences. However, one of the most preeminent algorithms in supervised learning is artificial neural network (ANN), inspired by the fundamental concept of neurons in the human brain and how they are interconnected [15]. These networks are based on the concept of neurons, which are basic units in the brain that communicate with each other to perform processes such as thinking and learning [16]. The algorithm simulates the functions of brain cells by proposing multiple layers of artificial neurons (an input layer and an output layer). These neurons interact with each other using weights assigned to each connection, and the role of the algorithm is to optimize these weights to minimize the error resulting from interactions with the input data, thereby producing accurate outputs [17]. Moreover, an older algorithm inspired by mathematics is the logistic regression (LR), which aims to find a perfect line that best fits the data points, minimizing the error between actual and predicted labels. These methods were used in statistical analyses before being adopted in ML [18]. The complexity of linear operations increased, leading to more sophisticated methods, such as support vector machine (SVM), where the main idea is to create clear boundaries between different data classes by maximizing the margin between them [19]. Comprehensively, most of ML classifiers have drawn their inspirations from mathematical operations or nature (e.g., simulating the functioning of human brain cells) to create robust systems (classifiers) for solving complex problems. Current ML classifiers face multiple challenges related to performance, accuracy or loss, overfitting, and handling data with complex and nonlinear patterns [20, 21].
@@
The proposed ALC has been evaluated using a variety of commonly used ML datasets, including Wine, Breast Cancer Wisconsin, Iris Flower, MNIST, and Voice Gender [24], which are explained in detail in Section 4.1. This diversity in the datasets ensures extensive coverage of different data types, including text, images, and audio, and enables handling binary and multi-class classification problems [25, 26, 27]. The purpose of using these datasets is to conduct comprehensive tests to assess the performance of the proposed ALC and compare it with the established classifiers. The originality and contributions that distinguish this research are as follows:
@@
This paper is structured as follows: Section 2 reviews the literature that has attempted to address classification issues across various data types. Section 3 provides an analytical overview of the human liver, focusing on detoxification function and the study’s motivation. Section 4 present the used materials and the proposed methodology, including the improvement of classifier design and FOX training algorithm. Sections 5 and 6 cover the presentation and analysis of results, including comparisons with previous works. Finally, the study concludes with findings, recommendations, limitations, and future research directions in Section 7.
@@ 3 Detoxification in Liver and Motivation
The liver, as illustrated in Figure 1, is the largest internal organ in the human body and is vital in numerous complex physiological processes. It is located in the right upper quadrant of the abdominal cavity and consists of two primary lobes, the right and left, surrounded by a thin membrane known as the hepatic capsule [36]. Internally, the liver is composed of microscopic units called hepatic lobules. These hexagonal structures contain hepatic cells organized around a central vein. The lobules are permeated by a network of hepatic sinusoids, which are small channels through which blood flows, facilitating the exchange of oxygen and nutrients between the blood and hepatic cells [37]. Furthermore, the liver receives blood from two sources, each contributing different functions. The oxygenated blood enters via the hepatic artery from the aorta, meeting the liver’s energy demands. While, the portal vein delivers nutrient-rich and toxin-rich blood from the gastrointestinal tract and spleen [38]. The blood from both sources mixes in the hepatic sinusoids, allowing the hepatic cells to perform metabolic and regulatory functions efficiently [39].
@@ 4.1 Materials
The following datasets are widely used by ML researchers to evaluate their work, making these benchmark datasets suitable for this paper. The MNIST dataset comprises 70,000 grayscale images of handwritten digits (0–9), each of size $28\times 28$ pixels. It is widely used for multi-class classification tasks due to its diversity and large size [45]. As this dataset is image-based, feature vectors were extracted using the linear discriminant analysis (LDA) method, which effectively reduces data dimensionality while emphasizing significant features. This process yielded nine features per image to construct the feature vectors [46]. Additionally, the Iris dataset, a small-scale collection containing 150 instances across three classes with four features per instance, was included in the proposed ALC evaluation [47, 48, 49]. The Breast Cancer Wisconsin dataset, a binary dataset containing 569 samples with 30 features each, was employed to assess the proposed ALC’s performance on high-dimensional data [50, 33]. Furthermore, the Wine dataset, consisting of 178 samples across three classes with 13 features per instance, was selected for its multi-class nature [48, 51]. Finally, the Voice Gender dataset was employed to ensure feature diversity. This dataset comprises 3,168 samples, each defined by 21 acoustic features, aimed at distinguishing gender (male or female) by leveraging unique vocal characteristics [35]. These datasets collectively provided a diverse range of classification challenges, enabling a comprehensive evaluation of the proposed ALC’s performance.
@@ 4.2.1 Artificial Liver Classifier
As explained earlier in Section 3, the detoxification process involves the liver’s ability to process toxins. Oxygenated blood enters the liver via the hepatic artery, while nutrient-rich blood flows through the portal vein. These sources mix within the hepatic sinusoids, enabling hepatic cells to perform essential functions, including a detoxification function that comprises two phases.
@@
$A_{ji}=\frac{1}{n}\sum_{k=1}^{n}(X_{jk}\times\text{cofactor}_{ki})+\frac{1}{fp}\sum_{k=1}^{f}\sum_{l=1}^{p}\text{cofactor}_{kl}$ (1)
where $A_{ji}$ is the matrix of reactive toxins, $X$ is the input toxins matrix (feature vector) and $n$ is the number of outputs (labels). The cofactor matrix is is initialized randomly within the range $[-1,1]$ and has dimensions $(f,p)$, where $f$ corresponds to the number of features in the input feature vector, and $p$ is the number of lobules. The human liver containing many lobules, with approximately 100,000 [52]. The $p$ play a crucial role in the phase I and should be tuned appropriately based on the problem. Thus, selecting $p$ with the range $100,000>p\geq f$ is typically sufficient and can often be determined through trial and error. Additionally, the term $\frac{1}{fp}\sum_{k=1}^{f}\sum_{l=1}^{p}\text{cofactor}_{kl}$ represents the mean of all elements in the cofactor matrix that used to balance the reaction.
However, the reactive toxins ($A$) must be activated to enhance their reactivity before progressing to phase II. This activation involves eliminating all negative values, effectively transforming them to zero while retaining only the positive values. This process is mathematically expressed by the following equation:
@@
where $A^{\prime}$ is the activated toxins matrix.
Phase II: involves the conjugation of modified compounds from phase I with water-soluble molecules to make them excretable. This phase reduces the toxicity of compounds and facilitates their elimination from the body. this phase can be mathematically modeled using Equation 1, but with key differences. Instead of toxins, the matrix $A^{\prime}$ is used as input, representing the modified compounds (activated toxins) generated in phase I. Additionally, a matrix referred to as the vitamin matrix is employed in place of the cofactor matrix. This vitamin matrix is initialized randomly within the range $[-1,1]$ and has dimensions $(p,n)$.
$B_{ji}=\frac{1}{n}\sum_{k=1}^{n}(A_{jk}\times\text{vitamin}_{ki})+\frac{1}{pn}\sum_{k=1}^{p}\sum_{l=1}^{n}\text{vitamin}_{kl}$ (3)
where $B_{ji}$ represents the conjugated compounds and $\frac{1}{pn}\sum_{k=1}^{p}\sum_{l=1}^{n}\text{vitamin}_{kl}$ represents the mean of all elements in the vitamin matrix.
Finally, after completing the reactions in both phase I and phase II, the detoxification process is concluded. The result is a set of less harmful and water-soluble wastes that can be excreted through bile, urine, stool, and other pathways. The elimination process is modeled using the softmax activation function, which provides probabilistic outputs for each class [53, 54, 55]. These probabilities represent the confidence scores for the likelihood of different detoxified compounds being excreted.
@@
The Algorithm 1 and Figure 2 describes the architecture of the proposed ALC. First, the cofactor and vitamin matrices are initialized randomly, where these matrices are defined based on the dimensions corresponding to the number of features ($f$), number of lobules ($p$), and number of output classes ($n$). Next, the IFOX, as presented in Algorithm 2, is configured, specifying the number of detoxification cycles (maximum number of epochs) and detoxification power (maximum number of fox agents). The target objective function for optimization is the REACTION procedure in Algorithm 1, where the IFOX optimizes the cofactor and vitamin matrices by minimizing the reaction error (i.e., loss). The optimized cofactor and vitamin matrices, resulting from the optimization process, are subsequently applied within the REACTION along with the toxins (feature vector) to predict the output classes.
Figure 2: Architecture of the proposed ALC.
@@
1:Input: toxins, number of features, number of lobules, number of outputs, detoxification cycles, and detoxification power.
@@
3:Initialize cofactor and vitamin matrices randomly.
4:Initialize the training algorithm (IFOX). $\triangleright$ Algorithm 2
5:Optimize cofactor and vitamin matrices using IFOX.
6:Reaction(toxins, optimized cofactor, optimized vitamin) to compute final predictions.
7:return predicted classes.
8:procedure Reaction(toxins, cofactor, vitamin)
9:     Compute the reactive toxins. $\triangleright$ using Equation 1
10:     Activate reactive toxins. $\triangleright$ using Equation 2
11:     Perform conjugation to make toxins less harmful. $\triangleright$ using Equation 3
12:     Normalize outputs to obtain predicted classes. $\triangleright$ using Equation 4
13:     return predicted classes.
14:end procedure
Furthermore, the flowchart visualized the proposed ALC is presented in Figure 3. Additionally, the source code for the implementation of the proposed ALC can be accessed at the following repository: https://github.com/mwdx93/alc.
@@ 4.2.2 Training Algorithm
    where $dim$ is the problem dimension, $Mint$ is the minimum time iteratively updated based on $T_{i}$, $a$ is an adjustment parameter computed as: $2\times(it-(\frac{1}{Max_{it}}))$, and $it$ is the current iteration.
@@
1:Input: Maximum number of epochs $epochs$, maximum number of fox agents $max_{fa}$
2:Output: $BestX$ and $BestFitness$
@@
The IFOX, as visualized in Algorithm 2, incorporates several improvements over the FOX. First, it transforms the balance between exploitation and exploration into a dynamic process using the $\epsilon$-greedy method, rather than a static approach [58, 59]. This dynamic adjustment is controlled by the parameter $\alpha$, which decreases progressively as the optimization process iterate. Second, the computation of distances is eliminated in favor of directly using the best position, facilitated by the parameter $\beta$, derived from $\alpha$. This modification simplifies the FOX by removing Equations 5 and 6, and simplifying Equation 8 by eliminating the probability parameter $p$ and the directional variables ($c_{1}$ and $c_{2}$). Third, in Equation 9, the variables $a$ and $Mint$ are excluded. Finally, the results shown in Figure 4 demonstrate that IFOX outperforms the FOX on the CEC2019 benchmark test functions, establishing it as a suitable training algorithm to integrate it with the proposed ALC.

### v2

@@ Abstract
Supervised machine learning classifiers sometimes face challenges related to the performance, accuracy, or overfitting. This paper introduces the Artificial Liver Classifier (ALC), a novel supervised learning model inspired by the human liver’s detoxification function. The ALC is characterized by its simplicity, speed, capability to reduce overfitting, and effectiveness in addressing multi-class classification problems through straightforward mathematical operations. To optimize the ALC’s parameters, an improved FOX optimization algorithm (IFOX) is employed during training. We evaluate the proposed ALC on five benchmark datasets: Iris Flower, Breast Cancer Wisconsin, Wine, Voice Gender, and MNIST. The results demonstrate competitive performance, with ALC achieving up to 100% accuracy on the Iris dataset–surpassing logistic regression, multilayer perceptron, and support vector machine–and 99.12% accuracy on the Breast Cancer dataset, outperforming XGBoost and logistic regression. Across all datasets, ALC consistently shows smaller generalization gaps and lower loss values compared to conventional classifiers. These findings highlight the potential of biologically inspired models to develop efficient machine learning classifiers and open new avenues for innovation in the field.
@@ 1 Introduction
In the early stages of AI, researchers focused on building systems (with minimal intelligence) capable of performing specific tasks using fixed rules (conditional and logical operations). As the field evolved, scientists realized that intelligent systems needed methods to learn from data, rather than relying on rigid rule-based methods with minimal capabilities [13, 14]. As a result, supervised learning algorithms, specifically classifiers, emerged as tools for learning systems to make predictions or decisions based on the available experiences. However, one of the most preeminent algorithms in supervised learning is artificial neural network (ANN), inspired by the fundamental concept of neurons in the human brain and how they are interconnected [15, 16]. These networks are based on the concept of neurons, which are basic units in the brain that communicate with each other to perform processes such as thinking and learning [17]. The algorithm simulates the functions of brain cells by proposing multiple layers of artificial neurons (an input layer and an output layer). These neurons interact with each other using weights assigned to each connection, and the role of the algorithm is to optimize these weights to minimize the error resulting from interactions with the input data, thereby producing accurate outputs [18]. Moreover, an older algorithm inspired by mathematics is the logistic regression (LR), which aims to find a perfect line that best fits the data points, minimizing the error between actual and predicted labels. These methods were used in statistical analyses before being adopted in ML [19]. The complexity of linear operations increased, leading to more sophisticated methods, such as support vector machine (SVM), where the main idea is to create clear boundaries between different data classes by maximizing the margin between them [20]. Comprehensively, most of ML classifiers have drawn their inspirations from mathematical operations or nature (e.g., simulating the functioning of human brain cells) to create robust systems (classifiers) for solving complex problems. Current ML classifiers face multiple challenges related to performance, accuracy or loss, overfitting, and handling data with complex and nonlinear patterns [21, 22].
@@
The proposed ALC has been evaluated using a variety of commonly used ML datasets, including Wine, Breast Cancer Wisconsin, Iris Flower, MNIST, and Voice Gender [25], which are explained in detail in Section˜4.1. This diversity in the datasets ensures extensive coverage of different data types, including text, images, and audio, and enables handling binary and multi-class classification problems [26, 27, 28]. The purpose of using these datasets is to conduct comprehensive tests to assess the performance of the proposed ALC and compare it with the established classifiers. The originality and contributions that distinguish this research are as follows:
@@
This paper is structured as follows: Section˜2 reviews the literature that has attempted to address classification issues across various data types. Section˜3 provides an analytical overview of the human liver, focusing on detoxification function and the study’s motivation. Section˜4 present the used materials and the proposed methodology, including the improvement of classifier design and FOX training algorithm. Sections˜5 and 6 cover the presentation and analysis of results, including comparisons with previous works. Finally, the study concludes with findings, recommendations, limitations, and future research directions in Section˜7.
@@ 3 Detoxification in Liver and Motivation
The liver, as illustrated in Figure˜1, is the largest internal organ in the human body and is vital in numerous complex physiological processes. It is located in the right upper quadrant of the abdominal cavity and consists of two primary lobes, the right and left, surrounded by a thin membrane known as the hepatic capsule [37]. Internally, the liver is composed of microscopic units called hepatic lobules. These hexagonal structures contain hepatic cells organized around a central vein. The lobules are permeated by a network of hepatic sinusoids, which are small channels through which blood flows, facilitating the exchange of oxygen and nutrients between the blood and hepatic cells [38]. Furthermore, the liver receives blood from two sources, each contributing different functions. The oxygenated blood enters via the hepatic artery from the aorta, meeting the liver’s energy demands. While, the portal vein delivers nutrient-rich and toxin-rich blood from the gastrointestinal tract and spleen [39]. The blood from both sources mixes in the hepatic sinusoids, allowing the hepatic cells to perform metabolic and regulatory functions efficiently [40].
@@ 4.1 Materials
The following datasets are widely used by ML researchers to evaluate their work, making these benchmark datasets suitable for this paper. The MNIST dataset comprises 70,000 grayscale images of handwritten digits (0–9), each of size $28\times 28$ pixels. It is widely used for multi-class classification tasks due to its diversity and large size [46]. To utilize the MNIST dataset with the proposed ALC, each image was preprocessed by flattening it to a vector of 784 dimensions. Each pixel was normalized, with its value transformed to have zero mean and unit variance to ensure consistent scaling. This was then followed by the use of linear discriminant analysis (LDA) to project or reduce the data into a low-dimensional space. Hence, LDA reduces each image to a nine-dimensional feature vector to effectively capture the most discriminative features while also reducing computational requirements [47]. Additionally, the Iris dataset, a small-scale collection containing 150 instances across three classes with four features per instance, was included in the proposed ALC evaluation [48, 49, 50]. The Breast Cancer Wisconsin dataset, a binary dataset containing 569 samples with 30 features each, was employed to assess the proposed ALC’s performance on high-dimensional data [51, 34]. Furthermore, the Wine dataset, consisting of 178 samples across three classes with 13 features per instance, was selected for its multi-class nature [49, 52]. Finally, the Voice Gender dataset was employed to ensure feature diversity. This dataset comprises 3,168 samples, each defined by 21 acoustic features, aimed at distinguishing gender (male or female) by leveraging unique vocal characteristics [36]. These datasets collectively provided a diverse range of classification challenges, enabling a comprehensive evaluation of the proposed ALC’s performance.
@@ 4.2.1 Artificial Liver Classifier
As explained earlier in Section˜3, the detoxification process involves the liver’s ability to process toxins. Oxygenated blood enters the liver via the hepatic artery, while nutrient-rich blood flows through the portal vein. These sources mix within the hepatic sinusoids, enabling hepatic cells to perform essential functions, including a detoxification function that comprises two phases.
In order to model the detoxification process in the liver, a biologically inspired computational model was chosen to be implemented, where every mathematical operation is merely treated as a simplistic representation of various known mechanisms regarding hepatic detoxification [41]. The major biochemical steps taken by hepatocytes in the elimination process are considered to include detoxification in two phases: oxidation (Phase I), which is carried out by specialized enzymes and cofactors, and conjugation (Phase II), which is also supported by specific enzymes and cofactors [43, 42]. In this formulation, these steps are coordinated by being transformed into operations in matrices. A set of molecular input toxins is first linearly overlaid by a cofactor matrix $C$ to model the oxidation action of cytochrome P450, followed by being passed through a nonlinear activation grid to model metabolite selection at a threshold. The second transformation is modeled in the form of conjugation through the interaction of vitamins, and this is subsequently normalized through the use of softmax to represent the classification of detoxified products. Although these operations are not intended to accurately reproduce the actual biochemical kinetics, they are carefully selected so that the structural and functional analogies are preserved—allowing the multi-staged, enzyme-driven, and spatially distributed nature of detoxification to be reproduced within a mathematically consistent and learnable system.
@@
$A_{ji}=\frac{1}{n}\sum_{k=1}^{n}(X_{jk}\cdot C_{ki})+\frac{1}{fp}\sum_{k=1}^{f}\sum_{l=1}^{p}C_{kl}$ (1)
where $A_{ji}$ is the matrix of reactive toxins, $X$ is the input toxins matrix and $n$ is the number of inputs. The $C$ is initialized randomly within the range [1,1][-1,1] and has dimensions $(f,p)$, where $f$ corresponds to the number of features in the input feature vector, and $p$ is the number of lobules. The term $\frac{1}{fp}\sum_{k=1}^{f}\sum_{l=1}^{p}C_{kl}$ represents the mean of all elements in the cofactor matrix $C$ that used to balance the reaction.
A human liver has a very large number of microscopic functional unit called lobules which are estimated to be of around 100,000 [53]. In our model the parameter $p$ is an abstraction of a range of choice in lobular diversity, and the columns of the cofactor matrix $C$ implement the various simulated lobular processing units. This structure allows introducing a spatial heterogeneity and an enzymatic variation, which can be witnessed in hepatic tissue, as part of the model. Although the biological premise of $p$ can be that of a number of lobules, direct insertion of a figure like $p=100,000$ in computation would be inefficient and also inconvenient because it has a high dimensionality and costs more computations. Thus, we introduce a practical range $f\leq p<100,000$. This will ensure there is diversity, and that every input feature interacts with a number of simulated lobules ensuring the model is computationally straightforward. The parameter $p$ in this range may be chosen by empirical methods of hyperparameter tuning, a compromise between the richness of the representation, and efficiency. It follows that the range is not arbitrary, but rather biological based on biological modeling under constrains. However, the reactive toxins ($A$) must be activated to enhance their reactivity before progressing to phase II. This activation involves eliminating all negative values, effectively transforming them to zero while retaining only the positive values. This process is mathematically expressed by the following equation:
@@
where $A^{\prime}$ is the activated toxins matrix. However, Equation 2 uses ReLU activation to imitate the biological selectivity that occurs in Phase I detoxification in which only reactive (positive) products pass to the next stage. This selection eliminates the non-reactive outputs keeps the computational efficiency and provides the non-linearity necessary in the downstream processing.
Phase II: involves the conjugation of modified compounds from phase I with water-soluble molecules to make them excretable. This phase reduces the toxicity of compounds and facilitates their elimination from the body. this phase can be mathematically modeled using Equation˜1, but with key differences. Instead of toxins, the matrix $A^{\prime}$ is used as input, representing the modified compounds (activated toxins) generated in phase I. Additionally, a matrix referred to as the vitamin matrix $V$ is employed in place of the cofactor matrix $C$. This matrix is initialized randomly within the range [1,1][-1,1] and has dimensions $(p,n)$.
$B_{ji}=\frac{1}{n}\sum_{k=1}^{n}(A_{jk}\cdot V_{ki})+\frac{1}{pn}\sum_{k=1}^{p}\sum_{l=1}^{n}V_{kl}$ (3)
where $B_{ji}$ represents the conjugated compounds and $\frac{1}{pn}\sum_{k=1}^{p}\sum_{l=1}^{n}V_{kl}$ represents the mean of all elements in the vitamin matrix $V$.
Lastly, when the reactions in Phase I and Phase II are finished, detoxification is then complete. The outcome is some less dangerous and water-soluble wastes that can be removed by means of bile, urine, stool, etc. The softmax activation function is used as a model of the elimination process: not only does it allow formulating a probabilistic output for each of the classes, but also captures the selective and competitive characteristic of biological excretion [54, 55, 56]. Many detoxified compounds simultaneously compete to be eliminated in the liver depending on issues such as solvency, the availability of transporters as well as priorities at the cellular level. Softmax reflects this behavior by placing higher probabilities on those compounds that are most dominant, or easily excreted, and thereby simulates the preferential clearance mechanism of the body.
@@
The  Algorithm˜1 and Figure˜2 describes the architecture of the proposed ALC. First, the cofactor and vitamin matrices are initialized randomly, where these matrices are defined based on the dimensions corresponding to the number of features ($f$), number of lobules ($p$), and number of output classes ($n$). Next, the IFOX, as presented in  Algorithm˜2, is configured, specifying the number of detoxification cycles (maximum number of training epochs) and detoxification power (maximum number of fox agents). The IFOX then optimizes the cofactor and vitamin matrices by minimizing the reaction error (i.e., loss). Finally, the optimized cofactor and vitamin matrices, resulting from this process, are subsequently used together with the toxin inputs (feature vectors) to predict the output classes.
@@
1:Input: toxins, number of features, number of lobules $p$, number of outputs, detoxification cycles, and detoxification power.
@@
3:Initialize cofactor matrix $C$ and vitamin matrix $V$ randomly.
4:Initialize the IFOX training algorithm. $\triangleright$ See Algorithm˜2
5:Optimize $C$ and $V$ using IFOX over defined cycles.
6:Compute reactive toxins. $\triangleright$ using Equation˜1
7:Activate reactive toxins. $\triangleright$ Phase I, using Equation˜2
8:Perform conjugation to make toxins less harmful. $\triangleright$ Phase II, using Equation˜3
9:Normalize outputs to obtain predicted classes. $\triangleright$ Elimination, using Equation˜4
10:return predicted classes $\hat{y}$.
Furthermore, the flowchart visualized the proposed ALC is presented in Figure˜3. Additionally, the source code for the implementation of the proposed ALC can be accessed at the following repository: https://github.com/mwdx93/alc, which includes the main ALC implementation, training scripts, and example datasets.
Figure 2: Architecture illustrating Phase I and Phase II reactions simulated by the proposed ALC, designed to mimic liver detoxification pathways.
@@ 4.2.2 Training Algorithm
    where dimdim is the problem dimension, $Mint$ is the minimum time iteratively updated based on $T_{i}$, $a$ is an adjustment parameter computed as: $2\times(it-(\frac{1}{Max_{it}}))$, and $it$ is the current iteration.
@@
1:Input: Maximum number of epochs epochsepochs, maximum number of fox agents $max_{fa}$
2:Output: BestXBestX and BestFitnessBestFitness
@@
The IFOX, as visualized in  Algorithm˜2, incorporates several improvements over the FOX. First, it transforms the balance between exploitation and exploration into a dynamic process using the $\epsilon$-greedy method, rather than a static approach [59, 60]. This dynamic adjustment is controlled by the parameter $\alpha$, which decreases progressively as the optimization process iterate. Second, the computation of distances is eliminated in favor of directly using the best position, facilitated by the parameter $\beta$, derived from $\alpha$. This modification simplifies the FOX by removing Equations 5 and 6, and simplifying Equation˜8 by eliminating the probability parameter $p$ and the directional variables ($c_{1}$ and $c_{2}$). Third, in Equation˜9, the variables $a$ and $Mint$ are excluded.